how do i Write y = –0.3x + 0.6 in standard form using integers.
Answer by TimothyLamb(4379)

---
given:
y = –0.3x + 0.6
---
standard form, decimal:
–0.3x - y = -0.6
0.3x + y = 0.6
---
standard form, fraction:
(3/10)x + y = (6/10)
---
standard form, integers:
3x + 10y = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php